OpenStudy (katiekruppa):

Caleb deposited money into a savings account each week. The table shows the balance in his account after different numbers of weeks. Caleb graphed this information. What is the slope of the line that represents this situation? Number of weeks 3 6 9 15 Account balance ($) 135 270 405 675

OpenStudy (katiekruppa):

help plz

OpenStudy (katiekruppa):

help plz

OpenStudy (anonymous):

I will help if you still ned it :D

OpenStudy (anonymous):

*need it XD

OpenStudy (anonymous):

slope = change in balance / change in weeks = (675 - 135)/(15 - 3) = 540/12 = 45